# Batch-ICL: Effective, Efficient, and Order-Agnostic In-Context Learning

### v1

@@ Abstract
In this paper, by treating in-context learning (ICL) as a meta-optimization process, we explain why LLMs are sensitive to the order of ICL examples. This understanding leads us to the development of Batch-ICL, an effective, efficient, and order-agnostic inference algorithm for ICL. Differing from the standard N-shot learning approach, Batch-ICL employs $N$ separate 1-shot forward computations and aggregates the resulting meta-gradients. These aggregated meta-gradients are then applied to a zero-shot learning to generate the final prediction. This batch processing approach renders the LLM agnostic to the order of ICL examples. Through extensive experiments and analysis, we demonstrate that Batch-ICL consistently outperforms most permutations of example sequences. In some cases, it even exceeds the performance of the optimal order for standard ICL, all while reducing the computational resources required. Furthermore, we develop a novel variant of Batch-ICL featuring multiple “epochs” of meta-optimization. This variant implicitly explores permutations of ICL examples, further enhancing ICL performance.
@@ 1 Introduction
Numerous studies (Olsson et al., 2022; Wang et al., 2023; Dai et al., 2023; Li et al., 2023; Akyürek et al., 2023; Von Oswald et al., 2023; Ren and Liu, 2023; Xie et al., 2022) have contributed to understanding the mechanism of ICL. Specifically, some research (Dai et al., 2023; Ren and Liu, 2023) describes ICL as a meta-optimization where an LLM is utilized as a meta-optimizer. Meta-gradients are produced through forward processing with ICL examples. These meta-gradients are then applied to the language model through the attention mechanism, resulting in an effective ICL model.
@@
We propose that these insights shed light on a commonly recognized issue in LLMs: their performance is highly influenced by the order of ICL examples. As emphasized by (Lu et al., 2022), changing the order of ICL examples can lead to significantly different outcomes. This paper offers a preliminary explanation for this phenomenon. In an $N$-shot ICL process, the meta-gradient is formed based on $N$ sequentially presented examples. Due to the causal attention mechanism in LLMs, the meta-gradient shaped by any given example is indirectly affected by the ones that came before it. Consequently, the order of these examples plays a critical role in shaping the final ICL model. This process is similar to training a neural network with $N$ samples where the batch size is one. In such cases, the gradients generated by each sample are influenced not only by the sample itself but also by the parameter updates from preceding samples, leading to variations in model parameters and overall performance. Given that the gradient from each sample represents a local optimum and is susceptible to causing suboptimal results, we suggest there is much potential for improving standard ICL performance by placing emphasis on the example order.
@@
In this paper, we introduce Batch-ICL, an effective, efficient and order-agnostic inference algorithm for in-context learning. Our method diverges from the standard ICL, which typically employs a single $N$-shot process for an $N$-shot ICL task. Instead, we implement $N$ separate 1-shot forward computations. This is then followed by the aggregation of $N$ corresponding meta-gradients at specific layers. These aggregated meta-gradients are subsequently employed in the forward computation when an LLM handling a zero-shot task, ultimately producing the final predictions.
@@
Batch-ICL addresses concerns related to the order of examples in ICL. Across a variety of tasks, Batch-ICL consistently exhibits improved accuracy compared to the average accuracy achieved through all permutations of ICL examples and occasionally outperforms even the best order permutation. Meanwhile, Batch-ICL reduces the computational resources needed for executing an ICL sequence.
@@
Additionally, we have discovered that although the standard ICL exhibits instability when presented with different ordered sequences, there are advantages to be gained from the interaction among sequential examples. We expand Batch-ICL into a “multi-epoch” variant, which implicitly enumerates the order permutations in a much more efficient manner, leading to further enhancements in performance.
The remainder of this paper is structured as follows: Section 2 offers an introduction to in-context learning viewed from a meta-optimization perspective. Section 3 presents our proposed Batch-ICL, including its algorithm pipeline, a complexity analysis, and multi-epoch extension. We demonstrate the experimental results and analysis in Section 4 and conduct extensive method analysis in Section 5. We discuss related works in Section 6 and conclude our paper in Section 7.
@@ 2.2 ICL is an Meta-Optimization
In an $N$-shot ICL setting, a Transformer (Vaswani et al., 2017)-based LLM consists of $K$ layers. The LLM processes an input sequence in the form of $I_{1},l_{1},I_{2},l_{2},\dots,I_{N},l_{N},I_{q}$, where $I_{i}$ and $l_{i}$ together represent a demonstration example, with $I_{i}$ being the input and $l_{i}$ as the corresponding label. Here, $I_{q}$ denotes the genuine query, and the LLM’s task is to predict its label. We represent the embedding of the input sequence at $k$-layer as $\boldsymbol{X}^{k}=[\boldsymbol{x}^{k}_{1},\boldsymbol{x}^{k}_{2},\dots,\boldsymbol{x}^{k}_{q}]$ with a special focus on $\boldsymbol{x}^{k}_{q}$, which represents the last query token’s representation. Without ambiguity, we will omit layer superscripts.
The output of an attention head can be expressed as follows:
$f(\boldsymbol{x}_{q})=\boldsymbol{W}_{V}\boldsymbol{X}\ \texttt{Softmax}\left(\frac{(\boldsymbol{W}_{K}\boldsymbol{X})^{T}\boldsymbol{W}_{Q}\boldsymbol{x}_{q}}{\sqrt{d_{out}}}\right),$ (4)
@@
3 Method
[…]
@@ 3.2 ICL Examples Batching and Meta-Gradients Aggregation
We adhere to the terminology used in Section 2. Considering a set of $N$ demonstration examples, we opt for the LLM to undertake $N$ individual 1-shot ICL learning processes, diverging from the standard single $N$-shot learning. Each sample received by LLMs is formatted as “$I_{i},l_{i},I_{q}$”, and $i$ spans the entire set $N$. At a selected layer $k$, we collect the $f_{i}(\boldsymbol{x_{q}})$ at the last position, i.e., the result of Eq. 5. Since $\boldsymbol{X}^{{}^{\prime}}$ in Eq. 5 now represents the representation of “$I_{i},l_{i}$”, as opposed to the entire set of ICL examples, we accordingly adjust the formula to: $f_{i}(\boldsymbol{x_{q}})=(\boldsymbol{W}_{0,i}+\Delta\boldsymbol{W}_{1,i})\boldsymbol{q}$.
@@
$\approx\left(\boldsymbol{W}_{0}+\frac{1}{N}\sum_{i=1}^{N}\Delta\boldsymbol{W}_{1,i}\right)q.$
@@ 3.3 Expanding to Multiple “Epochs”
We notice that the meta-optimization perspective can be extended to encompass “multiple epochs,” which in turn further improves the performance of Batch-ICL. To clarify this concept, let’s start by distinguishing the layers with superscripts $k$ in Eq. 5:
$f^{k+1}(\boldsymbol{x_{q}}^{k})$ $=f^{k+1}(f^{k}(\boldsymbol{x_{q}}^{k})).$ (7)
Here, we simplify the feedforward layer by treating it as a linear transformation, considering it as a unit matrix for clarity purposes. When we substitute the aggregated $\bar{f}^{k}(\boldsymbol{x_{q}}^{k})$ with its individual components $f^{k}_{i}(\boldsymbol{x_{q}}^{k})$, which are the separate outputs of each 1-shot ICL at layer $k$, and then derive the attention outputs from each $k+1$ layer for further aggregation, we actually engage in a form of meta-optimization during an additional epoch. Formally, the outputs of each 1-shot ICL at layer $k+1$ now turns to:
$f_{i}^{k+1}(\bar{f}^{k}(\boldsymbol{x_{q}}^{k}))$ (8)
$=(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\frac{1}{N}\sum_{j=1}^{N}\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$
$=(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})\frac{1}{N}\sum_{j=1}^{N}(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$
$=\frac{1}{N}\sum_{j=1}^{N}(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}.$
Notice the expression $(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$, which can be understood as conducting meta-optimization for example $j$ in the first $k$ layers, followed by a similar meta-optimization for example $i$ at layer $k+1$. We term this procedure an extra “epoch” within the Batch-ICL, and it can readily be extended to multi-epoch. Figure 1(c) illustrates this 2-epoch expansion of Batch-ICL.
@@ 4.1 Experimental Setup
$N$-shot Batch-ICL Best $N$-shot Batch-ICL Best
SST-2 $54.9_{\hskip 1.42271pt\text{[6.41]}}$   $\textbf{58.1}_{\hskip 1.42271pt\text{[8.78]}}$ $58.3_{\hskip 1.42271pt[8.89]}$ $63.8_{\hskip 1.42271pt[15.14]}$    $\textbf{68.8}_{\hskip 1.42271pt[11.35]}$    $69.4_{\hskip 1.42271pt[11.16]}$
RTE $53.0_{\hskip 1.42271pt\text{[6.77]}}$   $\textbf{62.9}_{\hskip 1.42271pt\text{[0.85]}}$ $64.2_{\hskip 1.42271pt[0.85]}$ $78.1_{\hskip 1.42271pt[4.21]}$    $\textbf{78.5}_{\hskip 1.42271pt[1.84]}$    $81.4_{\hskip 1.42271pt[2.30]}$
QNLI $50.7_{\hskip 1.42271pt\text{[1.79]}}$   $\textbf{52.8}_{\hskip 1.42271pt\text{[1.61]}}$ $55.3_{\hskip 1.42271pt[0.83]}$ $51.1_{\hskip 1.42271pt[2.51]}$    $\textbf{56.3}_{\hskip 1.42271pt[1.48]}$    $56.8_{\hskip 1.42271pt[1.14]}$
AGNews $28.7_{\hskip 1.42271pt\text{[3.84]}}$   $\textbf{29.5}_{\hskip 1.42271pt\text{[3.02]}}$ $29.6_{\hskip 1.42271pt[3.00]}$ $30.5_{\hskip 1.42271pt[6.28]}$    $\textbf{31.8}_{\hskip 1.42271pt[3.96]}$    $31.9_{\hskip 1.42271pt[3.80]}$
MRPC $40.5_{\hskip 1.42271pt\text{[10.5]}}$   $\textbf{66.7}_{\hskip 1.42271pt\text{[0.49]}}$ $67.1_{\hskip 1.42271pt[0.17]}$ $52.4_{\hskip 1.42271pt[14.83]}$    $\textbf{56.3}_{\hskip 1.42271pt[5.93]}$    $56.4_{\hskip 1.42271pt[5.95]}$
@@
Table 1: The comparison between Batch-ICL and standard $N$-shot ICL. We report the average score and standard deviation [$\sigma$] across 10 runs. The “Best” column reports the upper limit of Batch-ICL in which we search for the optimal $k$ for each test sample. Batch-ICL not only enhances LLMs’ ICL performance in diverse tasks but also decreases performance variability across different demonstration examples, often approaching its maximum potential in numerous scenarios. All results are passed with t-test with p-value < 0.05.
@@
The approach is evaluated on various classification tasks, containing one sentiment detection dataset SST-2 Socher et al. (2013), two natural language inference datasets RTE Dagan et al. (2006); Wang et al. (2019) and QNLI Wang et al. (2018), a topic classification dataset AGNews Zhang et al. (2015) and one paraphrase dataset MRPC Dolan and Brockett (2005).
@@
We evaluate performance on the test set when labels are accessible, and we sample 300 data from the validation set to derive the optimal value of $k$. In cases where test labels are not provided, we evaluate performance on the validation set and choose 300 data from the training set for obtaining $k$. In this comparison, we fix the value of $N$ at 4.
@@
For subsequent experiments in this paper, unless specifically stated otherwise, we present the average score obtained from 10 runs, each with different sampled demonstrations. This ensures reliable conclusions, as different ICL demonstrations lead to varying performance (Zhao et al., 2021; Perez et al., 2021).

### v3

@@ Abstract
In this paper, by treating in-context learning (ICL) as a meta-optimization process, we explain why LLMs are sensitive to the order of ICL examples. This understanding leads us to the development of Batch-ICL, an effective, efficient, and order-agnostic inference algorithm for ICL. Differing from the standard N-shot learning approach, Batch-ICL employs $N$ separate 1-shot forward computations and aggregates the resulting meta-gradients. These aggregated meta-gradients are then applied to the forward computation of a zero-shot query to generate the final prediction. This batch processing approach renders the LLM agnostic to the order of ICL examples. Through extensive experiments and analysis, we demonstrate that Batch-ICL consistently outperforms most permutations of ICL examples. In some cases, it even exceeds the performance of the best order for standard ICL, all while reducing the computational resources required. Furthermore, we develop a novel variant of Batch-ICL featuring multiple “epochs” of meta-optimization. This variant implicitly explores permutations of ICL examples, further enhancing ICL performance.111 Our code is available at https://github.com/Cardinalere/Batch-ICL.
@@ 1 Introduction
Numerous studies (Olsson et al., 2022; Wang et al., 2023b; Dai et al., 2023; Li et al., 2023; Akyürek et al., 2023; Von Oswald et al., 2023; Ren and Liu, 2023; Xie et al., 2022; Lv et al., 2024) have contributed to understanding the mechanism of ICL. Specifically, some research (Dai et al., 2023; Ren and Liu, 2023) describes ICL as a meta-optimization where an LLM is utilized as a meta-optimizer. Meta-gradients are produced through forward processing with ICL examples. These meta-gradients are then applied to the language model through the attention mechanism, resulting in an effective ICL model.
@@
We propose that these insights shed light on a commonly recognized issue: the ICL capacity of an LLM is highly influenced by the order of examples. As emphasized by (Lu et al., 2022), changing the order of ICL examples can lead to significantly different outcomes. This paper offers a preliminary explanation for this phenomenon: In an $N$-shot ICL process, the meta-gradient is formed based on $N$ sequentially presented examples. Due to the causal attention mechanism in LLMs, the meta-gradient shaped by any given example is indirectly affected by the ones that came before it. Consequently, the order of these examples plays a critical role in shaping the final ICL model. This process is similar to training a neural network with $N$ samples where the batch size is one. In such cases, the gradients generated by each sample are influenced not only by the sample itself but also by the parameter updated by preceding samples. The sample order leads to variations in parameters and the overall performance.
@@
Given that gradients from each sample represents a local optimum and is susceptible to causing suboptimal results, we suggest there is much potential for improving standard ICL. In this paper, we introduce Batch-ICL, an effective, efficient and order-agnostic inference algorithm for in-context learning. Our method diverges from the standard ICL, which typically employs a single $N$-shot process for an $N$-shot ICL task. Instead, we implement $N$ separate 1-shot forward computations. This is then followed by the aggregation of $N$ corresponding meta-gradients at a specific layer. These aggregated meta-gradients are then applied to the LLM in the same layer during the forward computation of a zero-shot query, ultimately generating the final predictions. This is equivalent to increase the meta-batch size in ICL from 1 to $N$, thereby reducing the randomness of meta-optimization and obtaining a better ICL model. Batch-ICL yields three key advantages:
@@
(1) Batch-ICL alleviates concerns related to the order of examples in ICL. Across a variety of tasks, Batch-ICL consistently exhibits improved accuracy compared to the average accuracy achieved through all permutations of ICL examples. It sometimes outperforms the best order. Meanwhile, Batch-ICL reduces the computational resources needed for executing an ICL sequence.
@@
(2) Additionally, we have discovered that although the standard ICL exhibits instability when presented with different ordered sequences, there are advantages to be gained from the interaction among sequential examples. We expand Batch-ICL into a “multi-epoch” variant, which implicitly enumerates the order permutations in a much more efficient manner, leading to further enhancements.
(3) In Batch-ICL, there is no limit to the number of demonstration examples provided that the length of each example does not surpass the pretrained context length. This effectively overcomes a significant constraint encountered in standard $N$-shot ICL.
@@ 3.1 ICL is an Meta-Optimization
In an $N$-shot ICL setting, a Transformer (Vaswani et al., 2017)-based LLM consists of $K$ layers. The LLM processes an input sequence in the form of $I_{1},l_{1},I_{2},l_{2},\dots,I_{N},l_{N},I_{q}$, where $I_{i}$ and $l_{i}$ together represent a demonstration example, with $I_{i}$ being the input and $l_{i}$ as the corresponding label. Here, $I_{q}$ denotes the genuine query, and the LLM’s task is to predict its label. We represent the embedding of the input sequence at $k$-layer as $\boldsymbol{X}^{k}=[\boldsymbol{x}^{k}_{1},\boldsymbol{x}^{k}_{2},\dots,\boldsymbol{x}^{k}_{q}]$ with a special focus on $\boldsymbol{x}^{k}_{q}$, which represents the last query token’s representation. Without ambiguity, we will omit layer superscripts. The output of an attention head can be expressed as follows:
$f(\boldsymbol{x}_{q})=\boldsymbol{W}_{V}\boldsymbol{X}\ \texttt{Softmax}\left(\frac{(\boldsymbol{W}_{K}\boldsymbol{X})^{T}\boldsymbol{W}_{Q}\boldsymbol{x}_{q}}{\sqrt{d_{out}}}\right),$
@@
In this section, we start by proposing an explanation (§$\S$3.2) about the sensitivity of LLMs to the order of ICL demonstration examples, which is built upon our comprehension of ICL as a meta-optimization process. Following this, in §$\S$3.3, we introduce our proposed solution, named “Batch-ICL,” in which the LLM performs an $N$-shot ICL by processing ICL examples in a “batch.” In §$\S$3.4, we elaborate on the extension of Batch-ICL to incorporate multiple “epochs,” thus fully harnessing its potential.
@@ 3.3 ICL Examples Batching and Meta-Gradients Aggregation
We adhere to the terminology used in Section 2. Considering a set of $N$ demonstration examples, we opt for the LLM to undertake $N$ individual 1-shot ICL learning processes, diverging from the standard single $N$-shot learning. Each sample received by LLMs is formatted as “$I_{i},l_{i},I_{q}$”, and $i$ spans the entire set $N$. At a selected layer $k$, we collect $f_{i}(\boldsymbol{x_{q}})$ at the last position, i.e., the result of Eq. 5. Since $\boldsymbol{X}^{{}^{\prime}}$ in Eq. 5 now represents the representation of “$I_{i},l_{i}$”, as opposed to the entire set of ICL examples, we accordingly adjust the formula to: $f_{i}(\boldsymbol{x_{q}})=(\boldsymbol{W}_{0,i}+\Delta\boldsymbol{W}_{1,i})\boldsymbol{q}$.
@@
$\approx(\boldsymbol{W}_{0}+\frac{1}{N}\sum_{i=1}^{N}\Delta\boldsymbol{W}_{1,i})\ \boldsymbol{q}.$
@@ 3.4 Expanding to Multiple “Epochs”
We notice that the meta-optimization perspective can be extended to encompass “multiple epochs,” which in turn further enhances Batch-ICL. To clarify this concept, let’s start by distinguishing the layers with superscripts $k$ in Eq. 5:
$f^{k+1}(\boldsymbol{x_{q}}^{k+1})$ $=f^{k+1}(f^{k}(\boldsymbol{x_{q}}^{k})).$ (7)
Here, we simplify the feed-forward layer by treating it as a linear transformation, considering it as a unit matrix for clarity purposes. When we substitute the aggregated $\bar{f}^{k}(\boldsymbol{x_{q}}^{k})$ for its individual components $f^{k}_{i}(\boldsymbol{x_{q}}^{k})$, which are the separate outputs of each 1-shot ICL at layer $k$, and then derive the attention outputs from each $k+1$ layer for further aggregation, we actually engage in a form of meta-optimization during an additional epoch. Formally, the outputs of each 1-shot ICL at layer $k+1$ now turns to:
[…]
Notice the expression $(\boldsymbol{W}_{0}^{k+1}+\Delta\boldsymbol{W}_{i}^{k+1})(\boldsymbol{W}_{0}^{k}+\Delta\boldsymbol{W}_{j}^{k})\boldsymbol{q}$, which can be understood as conducting meta-optimization for example $j$ in the first $k$ layers, followed by a similar meta-optimization for example $i$ at layer $k+1$. We term this procedure an extra “epoch” within the Batch-ICL, and it can readily be extended to multi-epoch. Figure 1(c) illustrates the 2-epoch expansion of Batch-ICL.
@@
3.5 Discussion on the Efficiency
We focus on the computational overhead related to the attention mechanism in Transformer models. This part is a critical efficiency bottleneck for Transformers, showcasing a time complexity of $O(n^{2})$ when processing texts of length $n$.
To analyze practical resource utilization, we tested the AGnews dataset utilizing $N$=4 on an A100-80G. Batch-ICL completes the inference of the entire dataset in 512.84 seconds, while the standard ICL method requires 697.09 seconds. Considering space efficiency, Batch-ICL provides $N$ times greater space efficiency than standard ICL under ideal conditions. In practice, executing the standard 4-shot AGnews task on Llama 2-13B consumes 32.1 GB of memory, while our method requires only 29.7 GB, wherein model parameters occupy 26 GB.
@@ 4.1 Experimental Setup
$N$-shot Batch-ICL Best PCW F-Ordered $N$-shot Batch-ICL Best PCW F-Ordered
SST-2 $54.9_{\hskip 1.42271pt\text{[6.4]}}$   $\textbf{58.7}_{\hskip 1.42271pt\text{[9.2]}}$ $58.8_{\hskip 1.42271pt[9.4]}$ $54.2_{\hskip 1.42271pt[6.9]}$ $56.4_{\hskip 1.42271pt[11.8]}$ $63.8_{\hskip 1.42271pt[15.1]}$    $\textbf{69.7}_{\hskip 1.42271pt[11.4]}$    $70.2_{\hskip 1.42271pt[11.5]}$ $51.9_{\hskip 1.42271pt[3.7]}$ $67.7_{\hskip 1.42271pt[10.9]}$
RTE $53.0_{\hskip 1.42271pt\text{[6.8]}}$   $\textbf{65.2}_{\hskip 1.42271pt\text{[1.9]}}$ $67.6_{\hskip 1.42271pt[0.5]}$ $56.6_{\hskip 1.42271pt[5.2]}$ $53.9_{\hskip 1.42271pt[5.1]}$ $78.1_{\hskip 1.42271pt[4.2]}$    $\textbf{79.2}_{\hskip 1.42271pt[3.5]}$    $80.9_{\hskip 1.42271pt[2.8]}$ $56.5_{\hskip 1.42271pt[7.3]}$ $80.5_{\hskip 1.42271pt[1.3]}$
QNLI $50.7_{\hskip 1.42271pt\text{[1.8]}}$   $\textbf{52.4}_{\hskip 1.42271pt\text{[1.3]}}$ $55.2_{\hskip 1.42271pt[0.7]}$ $50.1_{\hskip 1.42271pt[0.6]}$ $51.3_{\hskip 1.42271pt[0.8]}$ $51.1_{\hskip 1.42271pt[2.5]}$    $\textbf{56.7}_{\hskip 1.42271pt[0.9]}$    $56.9_{\hskip 1.42271pt[1.2]}$ $51.2_{\hskip 1.42271pt[1.1]}$ $56.2_{\hskip 1.42271pt[3.9]}$
AGNews $28.7_{\hskip 1.42271pt\text{[3.8]}}$   $29.3_{\hskip 1.42271pt\text{[3.1]}}$ $29.4_{\hskip 1.42271pt[3.0]}$ $25.8_{\hskip 1.42271pt[2.0]}$ $\textbf{30.2}_{\hskip 1.42271pt\text{[4.9]}}$ $30.5_{\hskip 1.42271pt[6.3]}$    $31.4_{\hskip 1.42271pt[3.9]}$    $31.6_{\hskip 1.42271pt[3.8]}$ $25.5_{\hskip 1.42271pt[0.8]}$ $\textbf{32.1}_{\hskip 1.42271pt[6.5]}$
MRPC $40.5_{\hskip 1.42271pt\text{[10.5]}}$   $\textbf{66.9}_{\hskip 1.42271pt\text{[0.5]}}$ $68.0_{\hskip 1.42271pt[0.2]}$ $56.8_{\hskip 1.42271pt[14.3]}$ $38.6_{\hskip 1.42271pt\text{[7.6]}}$ $52.4_{\hskip 1.42271pt[14.8]}$    $56.5_{\hskip 1.42271pt[0.6]}$    $67.9_{\hskip 1.42271pt[1.7]}$ $\textbf{63.8}_{\hskip 1.42271pt[8.0]}$ $63.3_{\hskip 1.42271pt[6.5]}$
@@
Table 1: Experimental results of classification tasks. We report the average score and standard deviation [$\sigma$] across 10 runs. The “Best” column reports the upper limit of Batch-ICL in which we search for the optimal $k$ for each test sample. We also compare Batch-ICL with PCW Ratner et al. (2023) and Fantastically Ordered Lu et al. (2022). Batch-ICL not only enhances LLMs’ ICL performance in diverse tasks but also decreases performance variability across different demonstration examples, often approaching its maximum potential. All bold results passed the t-test with a p-value < 0.05.
$N$-shot Batch-ICL
7B $66.38$ $65.91$
13B $67.88$ $67.63$
Table 2: BLEU scores for WMT2014 En-Fr.
@@
We evaluate Batch-ICL on classification tasks, containing one sentiment detection dataset SST-2 Socher et al. (2013), two natural language inference datasets RTE Dagan et al. (2006); Wang et al. (2019) and QNLI Wang et al. (2018), a topic classification dataset AGNews Zhang et al. (2015) and one paraphrase dataset MRPC Dolan and Brockett (2005). Moreover, we also utilize Batch-ICL for free-format generation tasks, such as machine translation. We choose the WMT2014 En-Fr (Bojar et al., 2014) benchmark.
Additionally, we compare Batch-ICL with Parallel Context Windows (PCW, Ratner et al., 2023) that also enhances the inference of ICL. In terms of parallel processed examples, PCW is the most related work to our study. PCW resets the position embedding for each example, implicitly handling these examples in parallel. However, it does not improve efficiency and introduces gaps in inference behavior, potentially compromising performance. We also compare Batch-ICL with Fantastically Ordered (F-Ordered, Lu et al., 2022), a prior study on selecting the optimal order for ICL. F-Ordered necessitates the enumeration of all permutations to identify an optimal prompt sequence.
@@
For classification tasks, we evaluate the accuracy on the test set when labels are accessible, and we sample 300 data from the validation set to derive the optimal value of $k$. In cases where test labels are not provided, we evaluate performance on the validation set and choose 300 data from the training set for obtaining $k$. For the machine translation task, we determine the optimal value of $k$ using the validation set and report the BLEU score Papineni et al. (2002).
@@
In this section, we fix the value of $N$ as 4. Unless specifically stated otherwise, we present the average score obtained from 10 runs, each with different sampled demonstration examples. This ensures reliable conclusions, as different ICL demonstrations lead to varying performance (Zhao et al., 2021; Perez et al., 2021).